Pls answer my question
When a ball is tossed to a height of 4 meters above the ground, it always rebounds to 40% of it's previous height until it stops. Find the total distance that the ball has covered when it strikes the ground for the fifth time.

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

each time the ball is tossed, it rebounds to 40% of the previous height.
when it is tossed, the height is 4 meters.
on the first rebound, the height is .4 * 4 = 1.6 meters.
on the second rebound, the height is .4 * 1.6 = .64 meters.
on the third rebound, the height is .4 * .64 = .256 meters.
etc.

at some point the height will decrease to 0.

this appears to be a geometric series.

the formula for the nth term of a geometric series is An = A1*r^(n-1).

the formula for the sum of the n terms of a geometric series is Sn = A1*(1-r^n)/(1-r).

if r is less than 1, than the infinite sum of the geometric series is S = A1/(1-r)

in your series, r is less than 1.
it is .4 to be exact.

the formula for the infinite sum of your series would be:
S = 4 / (1-r) = 4/.6 = 6.66666667 rounded to 8 decimal places.

to find the total distance traveled when the ball strikes the ground for the fifth time, the formula of Sn = A1*(1-r^n)/(1-r) is used.
that becomes:

S5 = 4*(1-.4^5)/.6 = 6.5984.

the ball will have traveled 6.5984 meters by the time it strikes the ground for the 5th time.

here's what the series looks like in excel.




Answer by ikleyn(52781)

.

            The solution by  @Theo is incorrect,  since he calculates the sums of distances in one direction at each bound,
            while the correct solution should account for them twice  (up and down).

            So,  I made the table in  Excel to provide an absolute clarity.

            See below.

 	4 m

-----------------------------
                up      down

1st rebound	1.6	1.6	
2nd rebound	0.64	0.64	
3rd rebound	0.256	0.256	
4th rebound	0.1024	0.1024	
5th rebound		 	
			
	total 	sum =	4 + the sum in the first column + the sum in the second column = 9.1968 meters.    ANSWER

Solved.